OpenStudy (anonymous):

find the area enclosed between the curve y=x^3 and the line y=x

OpenStudy (anonymous):

did you sketch the region?

OpenStudy (anonymous):

im not sure on how to

OpenStudy (anonymous):

you're not sure how to sketch x^3 ?

OpenStudy (anonymous):

no i mean i sketched it, i dont know what to do after

OpenStudy (anonymous):

say, for example, you're asking about the area enclosed for positive values of x..: 

OpenStudy (anonymous):

you can see that it's the area under the line x (from 0 to 1) minus the area under x^3 (from 0 to 1)

OpenStudy (anonymous):

aka:\[\int\limits_{0}^{1} x-x ^{3} dx\]

OpenStudy (anonymous):

can you evaluate that integral?

OpenStudy (anonymous):

and then how do we find the anti derivitve of that hat

OpenStudy (anonymous):

the usual way:\[\int\limits_{ }^{ } x^n dx = \frac{ x^{n+1} }{ n+1 }\]

OpenStudy (anonymous):

so for \[\int\limits_{ }^{} x^1 dx\] \[\int\limits_{ }^{} x^1 dx = \frac{ x^{1+1} }{ 1+1}\]

OpenStudy (anonymous):

try it for x^3...

OpenStudy (anonymous):

x^4/4

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

so \[\frac{ x^2 }{ 2 } -\frac{ x^4 }{4 } \]

OpenStudy (anonymous):

ok?

OpenStudy (anonymous):

and whats next?

OpenStudy (anonymous):

plug in the limts

OpenStudy (anonymous):

for the region on the interval x =0...+inf ( 1/2 -1/4 ) - (0/2 -0/4)

OpenStudy (anonymous):

kk